if n=10 and p=0.7 find the mean and standard deviation
Answer by stanbon(75887)

if n=10 and p=0.7 find the mean and standard deviation
---
mean = np = 10*0.7 = 7
=============
std = sqrt(npq) = sqrt(7*0.3) = sqrt(0.21) = 0.4583
====================
cheers,
Stan H.